# Robust Point Cloud Registration via Geometric Overlapping Guided Rotation Search

### v1

@@ V Stage I: Rotation Axis Search
where $\xi_{1}^{2}=2\sigma^{2}\chi_{1,0.95}^{2}$ is the 1D noise threshold.
The objective function (5) identifies the largest coplanar subset of $\{\mathbf{S}_{i}\}_{i=1}^{N}$ by fitting a plane with normal $\mathbf{r}$ at distance $d$ from the origin. Unlike TR-DE [27], we parameterize rotation axes via cube mapping [52] on the hemisphere, as shown in Fig. 3(a). The transformation from cube-face coordinates $(x_{f},y_{f})^{T}$ (e.g., XY-face) to Cartesian coordinate of rotation axis is
$\mathbf{r}_{\mathrm{p}}$ $=(x_{f},y_{f},1)^{T},$ (6)
@@
where $\mathbf{r}_{\mathrm{p}}$ is the unnormalized rotation axis $\mathbf{r}$ on the cube face. The initial search space per cube face is $1,1]\times[-1,1]$.
@@ V-A Projection onto Rotation Axis
The rotation axis $\mathbf{r}_{\mathrm{c}}$ for a surface element is represented by its center $\left((\underline{x}_{f}+\overline{x}_{f})/2,(\underline{y}_{f}+\overline{y}_{f})/2\right)^{T}$. The farthest possible axis $\mathbf{r}_{\mathrm{m}}$ from $\mathbf{r}_{\mathrm{c}}$ lies at the corner closest to the cube face center. The translation distance $d$ for correspondence $\mathbf{S}$ projected onto axis $\mathbf{r}$ is:
@@ V-B Gradually Converged Interval Stabbing
The boundaries of translation distance form $N$ intervals $\left\{[\underline{d}_{li},\overline{d}_{li}]\right\}_{i=1}^{N}$ and $\left\{[\underline{d}_{ui},\overline{d}_{ui}]\right\}_{i=1}^{N}$ respectively. The maximum accumulated inlier weights correspond to the peak interval overlap, as shown in Fig. 4(a) and (b). It can be estimated by interval stabbing algorithm [26] modified by weights: 1) Sort all interval boundaries by $d$, and assign the $i^{th}$ weight of left interval boundary to $w_{i}$, right boundary to $w_{i}$ for sweep line traversing from left to right, as shown in Fig. 4(c); 2) compute cumulative weights along $d$-axis, and identify the global maximum in Fig. 4(d).
@@
Figure 4: 1D RMQ solved by interval stabbing. (a) Difference vectors of correspondences projected onto rotation axis. (b) Estimated intervals $\underline{d},\overline{d}]$ and weights. (c) Sweep line for traversing discretized left interval boundary with $w_{i}$ weight and right with $w_{i}$ weight. (d) Accumulated weights and maximum overlapping.
@@ V-C Post Top-k Plane Fitting
Specifically, while the BnB search result for the rotation axis on the hemisphere yields the optimal solution for stage I, this axis may not correspond to the global optimum for the rigid transformation. We initialize the search by partitioning each cube face into four branches defined by corner coordinates $\left(0,0\right)^{T}$ and $\left(\pm 1,\pm 1\right)^{T}$. BnB search is then executed independently for each branch. The top-$k$ candidate rotation axes with highest $\underline{W}_{1}$ are selected as initial candidates for stage II.
@@
$\mathbf{\Gamma}_{S}$ $=\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\left(\mathbf{S}_{i}-\mathbf{S}_{c}\right)\left(\mathbf{S}_{i}-\mathbf{S}_{c}\right)^{T}=\mathbf{VDV}^{T},$ (13b)
@@ VI Stage II: Rotation Angle Search
where $\mathbf{p}_{i}$ and $\mathbf{q}_{i}$ denote the 2D projections of $\mathbf{P}_{i}$ and $\mathbf{Q}_{i}$ onto the plane perpendicular to the rotation axis, $\mathbf{R}_{\theta}$ represents the 2D rotation matrix parameterized by angle $\theta$, $\xi_{2}^{2}=2\sigma^{2}\chi_{2,0.95}^{2}$ defines the noise threshold based on the distribution of distances between two independent 2D points under Gaussian noise, and $\mathbf{c}$ is the rotation center.
@@ VI-A Rotation Center Determined by Angle
$\mathbf{c}(\theta)$ $=(x_{\mathbf{c}}(\theta),y_{\mathbf{c}}(\theta))^{T}$ (16)
@@
where $\mathbf{p}=(x_{p},y_{p})^{T}$ and $\mathbf{q}=(x_{q},y_{q})^{T}$ are 2D coordinates. Therefore, the range of rotation center $\mathbf{c}$ is determined by the range of angle $\underline{\theta},\overline{\theta}]$ and (16), as shown in Fig. 5(b). Given the prior Gaussian noise $\sigma$ in each dimension of correspondences, the covariance matrix of $\mathbf{c}$ is derived as:
@@
$=\left\|\xi_{\mathbf{c}}(\theta)\left(\mathbf{R}_{\frac{\theta}{2}}^{T}-\mathbf{R}_{\frac{\theta}{2}}\right)\mathbf{u}\right\|^{2}$ (21b)
@@
$=2\sigma^{2}\chi_{2,0.95}^{2}\left\|\mathbf{R}_{\frac{\pi}{2}}^{T}\mathbf{u}\right\|^{2}$ (21e)
@@ VI-B2 Segment tree with lazy propagation
Figure 6: Rotation center search using sweep line algorithm with segment tree. (a) Sweep line from left to right for traversing the rectangles. (b) Add a new segment to the tree and update the maximum overlapping within $O(\log N)$ time (e.g. a segment $y_{4},y_{8}]$ is added and only the $1\rightarrow 2\rightarrow 5\rightarrow y_{4}$ and $1\rightarrow 3$ paths are updated).
@@
where $\mathrm{SweepLine}(\cdot)$ computes the maximum rectangle overlap in Algorithm 2. The pruning strategy of BnB search follows stage I in Sec. V.
Algorithm 2 Sweep line algorithm with segment tree.
@@ VIII-1 Datasets and Metrics
$\mathrm{RE}$ $=\arccos\left(\frac{1}{2}\left(\mathrm{tr}\left(\hat{\mathbf{R}}^{T}\mathbf{R}\right)-1\right)\right)$ (29)
@@ VIII-2 Implementation details
We compare our method with the following baselines on real and synthetic datasets: 1) RANSAC [39] and its variant TCF [21], where RANSAC-1M and RANSAC-4M use a maximum of $10^{6}$ and $4\times 10^{6}$ iterations respectively; 2) graph-based methods: TEASER++ [43], SC2-PCR++ [46] and MAC [44]; 3) BnB-based methods: TEAR [28], HERE [29] and TR-DE [27]. Since TR-DE[27] also adopts a two-stage BnB search but is closed-source, we re-implement it for fair comparisons. The RANSAC implementation uses Open3D [60] in Python with multi-threading acceleration, and convergence parameters match the official FCGF [36] configuration. As SC2-PCR++ [46] is GPU-optimized whereas others are CPU-based, we adapt its source code for CPU execution following [29]. For multithreaded methods, the thread count is fixed at 12.
Moreover, implementations of RANSAC [39] in Open3D [60], TEASER++ [43], SC2-PCR++ [46], and MAC [44] provide their own feature matching methods. We use the default parameters to reproduce the results. As TEASER++ [43] provides cross checker to reject correspondences in feature space, we observed lower RR compared to configurations without this filter. Consequently, we disable this cross checker by default. For other methods, the features are matched by our preprocessing procedure in Section IV.
@@ VIII-A Indoor 3DMatch and 3DLoMatch Evaluation
Due to the large number of correspondences generated by FCGF [36], TEASER++ [43] requires over 2 hours to compute the maximum clique in some cases. Consequently, we enable its cross checker to reject most correspondences and display the corresponding time in gray in Table I SC2-PCR++ [46] benefits from its one-to-many feature matching, which increases the likelihood of resampling correct correspondences. This approach reduces FCGF’s [36] generalization error influence but incurs higher computational costs. Notably, SC2-PCR++ [46] achieves higher performance with learning-based descriptors than with traditional descriptors. This performance difference occurs because SC2-PCR++ [46] constructs feature graphs using normalized features, losing information from unnormalized FPFH [18] features during consensus set selection. Our method achieves suboptimal RR and the lowest RE and TE, with significantly reduced computation time compared to SC2-PCR++ [46] on identical hardware.
@@ VIII-B Outdoor KITTI Evaluation
Following [36], we evaluate compared methods using 555 pairs from KITTI test dataset (sequences 8 to 10). Traditional FPFH [18] and learning-based FCGF [36] descriptor are employed. Particularly, FCGF [36] provides normalized and non-normalized 32-dimensional pre-trained features at 30 cm resolution, and we report the results of normalized version. To ensure predictable execution time and manageable memory consumption, each point cloud is randomly downsampled to 8000 points for pairwise registration. Under the Atlanta World assumption [61] that is already applied to LiDAR point cloud [62], we constrain rotation axis search to the cube face with normal of z-axis, and set top-$k$ to 1. The preprocessing distance factor $d_{\mathrm{f}}$ of feature matching is set to 0.05 and the number of kNN searched features $k_{\mathrm{f}}$ is set to 10 for both FPFH [18] and FCGF [36] descriptors. The registration is successful if RE $5^{\circ}$ and TE <1<1 m.
@@
VIII-C Performance Evaluation on Synthetic Data
To validate the robustness, space and time complexity of our method versus SOTA methods, we evaluate them with varying number of correspondences and outlier ratio. The Armadillo model from the Stanford 3D Scanning Repository [55] serves as synthetic data. This model consists of 172974 points, and we downsample to $N\in\{1000,5000,\dots,100000\}$ correspondences. Point clouds are normalized to a $1,1]^{3}$ m cube before adding zero-mean Gaussian noise and outliers. For the outlier ratio, we randomly generate points in a 10 m diameter sphere, and replace the target points with a probability equal to outlier ratio. Noise magnitude and inlier threshold $\xi$ (set to 0.1) correspond to Eq. (1). We record the estimated RE and TE, elapsed time and memory usage of compared methods across number of correspondences and outlier ratios in synthetic experiments.
VIII-C1 Number of correspondences

### v2

@@ V Stage I: Rotation Axis Search
where $\xi_{1}^{2}=2\sigma^{2}\chi_{1,0.95}^{2}$ is the 1D noise threshold ($\chi_{1,0.95}\approx 1.96$).
The objective function (5) identifies the largest coplanar subset of $\{\mathbf{S}_{i}\}_{i=1}^{N}$ by fitting a plane with normal $\mathbf{r}$ at distance $d$ from the origin. Unlike TR-DE [13], we parameterize rotation axes via cube mapping [60] on the hemisphere, as shown in Fig. 3(a). The transformation from cube-face coordinates $(x_{f},y_{f})^{\top}$ (e.g., XY-face) to Cartesian coordinate of rotation axis is
$\mathbf{r}_{\mathrm{p}}$ $=(x_{f},y_{f},1)^{\top},$ (6)
@@
where $\mathbf{r}_{\mathrm{p}}$ is the unnormalized rotation axis $\mathbf{r}$ on the cube face. The initial search space per cube face is $1,1]\times[-1,1]$. Moreover, a comparative study of different spherical parameterizations (e.g. Miller projection used in TR-DE [13]) is presented in Appendix B.
@@ V-A Projection onto Rotation Axis
The rotation axis $\mathbf{r}_{\mathrm{c}}$ for a surface element is represented by its center $\left((\underline{x}_{f}+\overline{x}_{f})/2,(\underline{y}_{f}+\overline{y}_{f})/2\right)^{\top}$. The farthest possible axis $\mathbf{r}_{\mathrm{m}}$ from $\mathbf{r}_{\mathrm{c}}$ lies at the corner closest to the cube face center. The translation distance $d$ for correspondence $\mathbf{S}$ projected onto axis $\mathbf{r}$ is:
@@ V-B Gradually Converged Interval Stabbing
The boundaries of translation distance form $N$ intervals $\left\{[\underline{d}_{li},\overline{d}_{li}]\right\}_{i=1}^{N}$ and $\left\{[\underline{d}_{ui},\overline{d}_{ui}]\right\}_{i=1}^{N}$ respectively. The maximum accumulated inlier weights correspond to the peak interval overlap, as shown in Fig. 4(a) and (b). It can be estimated by interval stabbing algorithm [49] modified by weights: 1) Sort all interval boundaries by $d$, and assign the $i^{th}$ weight of left interval boundary to $w_{i}$, right boundary to $w_{i}$ for integrating from left to right, as shown in Fig. 4(c); 2) compute cumulative weights along $d$-axis, and identify the global maximum in Fig. 4(d).
@@
Figure 4: 1D RMQ solved by interval stabbing. (a) Difference vectors of correspondences projected onto rotation axis. (b) Estimated intervals $\underline{d},\overline{d}]$ and weights. (c) Integrating discretized left interval boundary with $w_{i}$ weight and right with $w_{i}$ weight. (d) Accumulated weights and maximum overlapping.
@@ V-C Post Top-k Plane Fitting
Specifically, while the BnB search result for the rotation axis on the hemisphere yields the optimal solution for stage I, this axis may not correspond to the global optimum for the rigid transformation. We initialize the search by partitioning each cube face into four branches defined by corner coordinates $\left(0,0\right)^{\top}$ and $\left(\pm 1,\pm 1\right)^{\top}$. BnB search is then executed independently for each branch. The top-$k$ candidate rotation axes with highest $\underline{W}_{1}$ are selected as initial candidates for stage II.
@@
$\mathbf{\Gamma}_{S}$ $=\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\left(\mathbf{S}_{i}-\mathbf{S}_{c}\right)\left(\mathbf{S}_{i}-\mathbf{S}_{c}\right)^{\top}=\mathbf{VDV}^{\top},$ (13b)
@@ VI Stage II: Rotation Angle Search
where $\mathbf{p}_{i}$ and $\mathbf{q}_{i}$ denote the 2D projections of $\mathbf{P}_{i}$ and $\mathbf{Q}_{i}$ onto the plane perpendicular to the rotation axis, $\mathbf{R}_{\theta}$ represents the 2D rotation matrix parameterized by angle $\theta$, $\xi_{2}^{2}=2\sigma^{2}\chi_{2,0.95}^{2}$ defines the noise threshold based on the distribution of distances between two independent 2D points under Gaussian noise ($\chi_{2,0.95}\approx 2.448$), and $\mathbf{c}$ is the rotation center.
@@ VI-A Rotation Center Determined by Angle
$\mathbf{c}(\theta)$ $=(x_{\mathbf{c}}(\theta),y_{\mathbf{c}}(\theta))^{\top}$ (16)
@@
where $\mathbf{p}=(x_{p},y_{p})^{\top}$ and $\mathbf{q}=(x_{q},y_{q})^{\top}$ are 2D coordinates. Therefore, the range of rotation center $\mathbf{c}$ is determined by the range of angle $\underline{\theta},\overline{\theta}]$ and (16), as shown in Fig. 5(b). Given the prior Gaussian noise $\sigma$ in each dimension of correspondences, the covariance matrix of $\mathbf{c}$ is derived as:
@@
$=\left\|\xi_{\mathbf{c}}(\theta)\left(\mathbf{R}_{\frac{\theta}{2}}^{\top}-\mathbf{R}_{\frac{\theta}{2}}\right)\mathbf{u}\right\|^{2}$ (21b)
@@
$=2\sigma^{2}\chi_{2,0.95}^{2}\left\|\mathbf{R}_{\frac{\pi}{2}}^{\top}\mathbf{u}\right\|^{2}$ (21e)
@@ VI-B2 Segment tree with lazy propagation
Figure 6: Rotation center search using sweep line algorithm with segment tree. (a) 2D RMQ with circular bounding circles. (b) 2D RMQ with axis-aligned rectangles can be solved in $O(N\log N)$ time. (c) Sweep line from left to right for traversing the rectangles. (d) Add a new segment to the tree and update the maximum overlapping within $O(\log N)$ time (e.g. a segment $y_{4},y_{8}]$ is added and only the $1\rightarrow 2\rightarrow 5\rightarrow y_{4}$ and $1\rightarrow 3$ paths are updated).
@@
where $\mathrm{SweepLine}(\cdot)$ computes the maximum rectangle overlap in Algorithm 2. The pruning strategy of BnB search follows stage I in Sec. V. The searching dimension in stage II is compressed to 1D by the proposed RMQ equivalency, as shown in Fig. 7.
Figure 7: Branching strategy for the 1D rotation angle interval $\underline{\theta},\overline{\theta}]$. Green lines denote inlier correspondences, while red lines indicate outliers. After branching of the current node, sub-branches with low upper bounds are pruned.
Algorithm 2 Solving 2D RMQ by sweep line algorithm.
@@ VI-D Post-Refinement with IRLS
VI-E Subproblem of 4-DoF Registration
According to Fig. 2, 4-DoF registration with a known rotation axis can be viewed as a constrained subproblem of the original 6-DoF rigid registration. Under this setting, the proposed framework admits two alternative decomposition strategies, depending on the order in which the remaining DoF are resolved:
  1. 1.
    1+3 DoF decomposition: The translation along the known rotation axis is firstly estimated via interval stabbing to filter out outliers, similar in spirit to the approach of Li et al. [40]. Given the estimated translation, the remaining 2D rigid transformation is then solved using the proposed method.
  2. 2.
    3+1 DoF decomposition: The 2D rigid transformation on the plane perpendicular to the rotation axis is firstly estimated, together with outlier filtering on the projected plane. The remaining 1D translation along the rotation axis is subsequently recovered via interval stabbing.
The two decomposition strategies exhibit different robustness characteristics: (1) The 1+3 DoF decomposition relies on the consistency of 1D projected correspondences and is therefore sensitive to outlier distributions along the axis; (2) The 3+1 DoF decomposition firstly constrains 3 DoF, leveraging stronger geometric structure in the projected plane, resulting in more informative residuals and a reduced ambiguity for the subsequent 1D translation estimation. This ordering is expected to provide improved robustness under high outlier ratios or noisy correspondences, which is further experimentally examined in Sec. VIII-C.
@@ VIII-1 Datasets and Metrics
$\mathrm{RE}$ $=\arccos\left(\frac{1}{2}\left(\mathrm{tr}\left(\hat{\mathbf{R}}^{\top}\mathbf{R}\right)-1\right)\right)$ (29)
@@ VIII-2 Implementation details
We compare our method with the following baselines on real and synthetic datasets: 1) RANSAC [24] and its variant TCF [58], where RANSAC-1M and RANSAC-4M use a maximum of $10^{6}$ and $4\times 10^{6}$ iterations respectively; 2) graph-based methods: TEASER++ [70], SC2-PCR++ [14], and MAC [78]; 3) BnB-based methods: TEAR [32], HERE [31], and TR-DE [13]. Since TR-DE[13] also adopts a two-stage BnB search but is closed-source, we re-implement it for fair comparisons. The RANSAC implementation uses Open3D [81] in Python with multi-threading acceleration, and convergence parameters match the official FCGF [17] configuration. As SC2-PCR++ [14] is GPU-optimized whereas others are CPU-based, we adapt its source code for CPU execution following [31]. For multithreaded methods, the thread count is fixed at 12.
Moreover, implementations of RANSAC [24] in Open3D [81], TEASER++ [70], SC2-PCR++ [14], and MAC [78] provide their own feature matching pipelines. We adopt the default parameter settings recommended by the respective authors to reproduce the reported results. TEASER++ [70] optionally employs a cross-checker (CC) in feature space to reject ambiguous correspondences. We observe that enabling this filter leads to a lower RR compared to configurations without CC. Accordingly, we report the results of TEASER++ [70] both without and with the CC, denoted as TEASER++ and TEASER++(w/ CC), respectively. For the remaining methods, feature correspondences are generated using the same preprocessing procedure described in Section IV to ensure a fair comparison.
For the 4-DoF pose estimation subproblem under gravity priors, we evaluate both the proposed 1+3 and 3+1 decomposition strategies. The comparisons include representative 4-DoF registration methods, namely BnB and FMP+BnB [10], Quatro [41], Li et al. [40], and the correspondence-free 3D-BBS [2]. All 4-DoF methods are evaluated using the same ground-truth rotation axis.
@@ VIII-A Indoor 3DMatch and 3DLoMatch Evaluation
Due to the large number of correspondences generated by FCGF [17], TEASER++ [70] requires over 2 hours and runs out-of-memory to compute the maximum clique in some cases. SC2-PCR++ [14] benefits from its one-to-many feature matching, which increases the likelihood of resampling correct correspondences. This approach reduces FCGF’s [17] generalization error influence but incurs higher computational costs. Notably, SC2-PCR++ [14] achieves higher performance with learning-based descriptors than with traditional descriptors. This performance difference occurs because SC2-PCR++ [14] constructs feature graphs using normalized features, losing information from unnormalized FPFH [55] features during consensus set selection. Our method achieves suboptimal RR and the lowest RE and TE, with significantly reduced computation time compared to SC2-PCR++ [14] on identical hardware.
@@ VIII-B Outdoor KITTI Evaluation
Following [17], we evaluate compared methods using 555 pairs from KITTI test dataset (sequences 8 to 10). Traditional FPFH [55] and learning-based FCGF [17] descriptor are employed. Particularly, FCGF [17] provides normalized and non-normalized 32-dimensional pre-trained features at 30 cm resolution, and we report the results of normalized version. To ensure predictable execution time and manageable memory consumption, each point cloud is randomly downsampled to 8000 points for pairwise registration. Under the Atlanta World assumption [61] that is already applied to LiDAR point cloud [42], we constrain rotation axis search to the cube face with normal of z-axis, and set top-$k$ to 1. The preprocessing distance factor $d_{\mathrm{f}}$ of feature matching is set to 0.05 and the number of kNN searched features $k_{\mathrm{f}}$ is set to 10 for both FPFH [55] and FCGF [17] descriptors. The registration is successful if RE $5^{\circ}$ and TE <2<2 m.
@@
VIII-C 4-DoF LiDAR Registration with Gravity Priors
Under gravity priors, LiDAR registration reduces the registration to 4 DoF with a known rotation axis. In this section, we evaluate the proposed method using both the 1+3 and 3+1 decomposition strategies, and compare it with representative 4-DoF registration methods on the KITTI dataset. To ensure a fair comparison, the ground-truth rotation axis is provided to all methods.
TABLE IV: 4-DoF registration results on KITTI Dataset.
FPFH + rotation axis prior
[…]
Quatro [41] 99.64 0.13 9.45 2.10
BnB [10] 98.92 0.19 16.54 0.17
FMP+BnB [10] 99.64 0.27 24.91 0.05
Li et al. [40] 94.59 0.24 28.01 0.62
3D-BBS [2] 99.82 0.62 101.06 8.38
GMOR (1+3) 99.64 0.08 7.80 0.05
GMOR (3+1) 100.00 0.08 7.18 0.17
Figure 12: Initial feature correspondences and registration result under the BEV view of LiDAR point clouds, where the source point cloud is shown in orange and the target in blue. (a) Initial feature correspondences, where inliers are shown as green lines and outliers as red lines; $\mathbf{c}$ denotes the ground-truth Chasles’ rotation center. (b) Registration result obtained by our method.
As shown in Table IV, the proposed method achieves competitive performance under the 4-DoF setting. Among the two decomposition strategies, the 3+1 variant consistently yields the most robust results, achieving the highest recall together with the lowest transformation error. In comparison, the 1+3 decomposition firstly estimates the translation along the rotation axis, remaining efficient but exhibits slightly higher sensitivity to outliers due to the 1D geometric constraint. As illustrated in Fig. 12, our method jointly enforces both rotation and translation on the 2D plane, enabling accurate inlier identification and yielding correct transformation. These results empirically validate the analysis in Sec. VI-E, where resolving the 2D rigid transformation firstly imposes stronger geometric constraints and reduces ambiguity for the subsequent 1D translation estimation.
@@
Compared with existing 4-DoF registration methods, the proposed method achieves strong robustness while retaining computational efficiency. Quatro [41] exhibits similar robustness with improved runtime efficiency compared to previous 6-DoF TEASER++ [70] on KITTI dataset. BnB and FMP+BnB [10] perform translation-space search with interval stabbing of rotation angle after ISS-based correspondence filtering, while FMP further suppresses outliers, the reduced number of retained correspondences leads to higher efficiency but lower registration accuracy. Li et al. [40] employ a three-stage pipeline combining interval stabbing, BnB searches and global voting, which is more sensitive to outlier distributions and results in lower recall. Finally, the correspondence-free 3D-BBS [2] achieves high recall at the cost of significantly increased runtime and TE, reflecting the trade-off between robustness and efficiency imposed by voxel map resolution.
VIII-D Performance Evaluation on Synthetic Data
To validate the robustness, space and time complexity of our method versus SOTA methods, we evaluate them with varying number of correspondences and outlier ratio. The Armadillo model from the Stanford 3D Scanning Repository [18] serves as synthetic data. This model consists of 172974 points, and we downsample to $N\in\{1000,5000,\dots,100000\}$ correspondences. Point clouds are normalized to a $1,1]^{3}$ m cube before adding zero-mean Gaussian noise and outliers. For the outlier ratio, we randomly generate points in a 10 m diameter sphere, and replace the target points with a probability equal to outlier ratio. Noise magnitude and inlier threshold $\xi$ (set to 0.1) correspond to (1). We record the estimated RE and TE, elapsed time and memory usage of compared methods across number of correspondences and outlier ratios in synthetic experiments.
VIII-D1 Number of correspondences